An open box is to be constructed from a square piece of sheet metal by removing a square of side 6 feet from each corner and turning up the edges. If the box is hold to 864 cubic feet, what should be the dimensions of the sheet metal?
Answer by checkley77(12844)

6(x-12)(x-12)=864
6(x^2-24x+144)=864
6x^2-144x+864=864
6x^2-144x=0
6x(x-24)=0
x-24=0
x=24 feet for the sides of the square piece of sgheet metal.
proof:
6(24-12)(24-12)=864
6*12*12=864
864=864